OpenStudy (anonymous):

Arithmetic Sequence question - What is the d value in the following arithmetic sequence? -7, -2, 3, 8, .. please help:)

hartnn (hartnn):

the common difference d is just the difference between the consecutive terms. d= 2nd term - 1st term = 3rd term - 2nd term = ...and so on

hartnn (hartnn):

here d = -2 - (-7) = ... or d = 3- (-2) = ... or d = 8-3 = ...

OpenStudy (anonymous):

so the D value is basically an equation?

hartnn (hartnn):

no, its a number its the difference between consecutive terms

hartnn (hartnn):

let me give you an example 10,20,30,40.... here d = 20-10 = 30-20 = 40-30 = \(\ \Large 10 \)

OpenStudy (anonymous):

So the D value for -7, -2, 3, 8, is 5?

hartnn (hartnn):

thats correct! :)

OpenStudy (anonymous):

thanks so much! <3 :)

hartnn (hartnn):

most welcome ^_^ <3